I am lost

Solve the system by substitution.
3x – y = –7
x + y = –9

Answer by scott8148(6628)

from first equation 3x+7=y ... using second equation x+(3x+7)=-9 ... 4x+7=-9 ... 4x=-16 ... x=-4

3(-4)-y=-7 ... -12+7=y ... -5=y